Calculate the required volume needed to make a 1.65 M solution consisting of 33.8 moles.

Calculate the required volume needed to make a 3.14 M solution consisting of 2.53 g of H2SO4

Calculate the number of moles of barium sulfide required to make 200 mL of a 0.53 M solution.

Calculate the mass of sodium chloride in a 0.0854 M solution, given 625 mL.

511 mL of a 9.22 M solution was diluted to 18.5 L. What is the new concentration?

A chemist prepared 112.8 L of a 0.0373 M solution using a 12.4 M stock. What volume of stock solution was needed?

A chemist combines 140. mL of a 1.25 M solution of sodium carbonate with 100. mL of a 0.750 M solution of silver nitrate. What is the mass of precipitate generated? What is the concentration of sodium nitrate generated?
Answer #1

Solution @ molarity = (mole)solute volume of solution 1165 M = (33.8) O solution Vsolution = 33.8 =( 20.48 litre) 165 ( molar0.0854 = (w+) plass X loco 58.5 X 625 (ct)noni = 0.0854 * 58.5x625 ovo = 3.122 gm As the solution is getting Diluted, so appl3 Nacou (.) +2 ag 103 (anyo Ag, costvarez [240 me, 1.25m) [come, 0.75m] [PP+] = 1:25 X 140 low 0.75 X 100 loco moles = 0.175concentration of 0.075 moles y love 240 me. Nanos = (0-3125 M) CS Scanned with CamScanner